OpenStudy (anonymous):

36a^(2x)-b^2 i need help i cant figure out the answer

OpenStudy (cwrw238):

this is the difference of 2 squares you need to find the square root of 36a^(2x) and b^2

OpenStudy (cwrw238):

thats the first part can you do that?

OpenStudy (anonymous):

i can try

OpenStudy (cwrw238):

what about b^2? whats square root of this?

OpenStudy (cwrw238):

b^2 = b*b so by definition square root of b^2 = b the square root of something is another something which when multiplied by itself gives the first something

OpenStudy (cwrw238):

eg ssquare root of 9 is 3 because 3*3 = 9 square root of b^4 = b^2 because b^2 * b^2 = b^4